Find the foci of an ellipse with the equation 7(x - 2)2 + 3(y - 2)2 = 21.
Answer by fractalier(6550)

Put it in standard form first...
7(x - 2)^2 + 3(y - 2)^2 = 21 we get
(x - 2)^2 / 3 + (y - 2)^2 / 7 = 1
the center of this ellipse is (2, 2)...it is taller than it is wide...
the distance from the center to the focus is found by
c^2 = b^2 - a^2 = 7 - 3 = 4 so that c = 2 and the foci are at
(2, 0) and (2, 4)